The perimeter of a rectangular window is 34 feet. If the width of the window is three feet more than the length, what are the dimensions of the window?
Answer by JBarnum(2146)

p=2(l+w)
w=3+l
34=2(l+(l+3))
17=2l+3
14=2l
7=l
10=w
check
10+10+7+7=34
34=34
correct